How do you solve for the area of a shaded region? I have an unshaded circle within a shaded circle. The formula I was given is pi r squared. The numbers given are in the two circles inside the shaded region each have a diameter of 3 and the shaded region is twelve. I tried squaring them and then subtracting. (it says to subtract the shaded region)
Answer by stanbon(75887)

How do you solve for the area of a shaded region? I have an unshaded circle within a shaded circle. The formula I was given is pi r squared. The numbers given are in the two circles inside the shaded region each have a diameter of 3 and the shaded region is twelve. I tried squaring them and then subtracting. (it says to subtract the shaded region)
-----
Find the area of the larger circle.
A = pi*12^2 = 144pi
-----
Find the area of the inner circle.
a = pi*3^2 = 9pi
----
Subtract the inner circle area from the larger circil area
to find the shaded area.
Area = 144pi-9pi = 135pi sq units.
===============
Cheers,
Stan H.
==============